RS04 postcal data collected after deployment on MOBY265 with the BS03.
Problems seen (and solved) and good things:
Track pixels: Track: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 Left: 1 74 150 223 296 369 440 512 585 656 729 802 875 951 Right: 73 149 222 295 368 439 511 584 655 728 801 875 950 1024
Wavelength Cals
P(2,:) = [-5.045113e-09 1.953268e-05 0.3859776 502.4691];
P(6,:) = [-5.491718e-09 2.044926e-05 0.3852714 502.6577];
P(13,:) = [-5.237887e-09 1.991368e-05 0.3857711 502.4651];
All the raw data files and their images , Table of KEYWORDS variable | |||
Page Number |
Link |
Description |
Date |
1.01 |
Day 01 - 16 Oct Raw Data |
DAY01 - led testing (light is not showing up on the right tracks) |
30 Nov 2018 |
1.02 |
Day 02 - 29 Nov Raw Data |
DAY02 - Es postcal sysrsp - Es rsp synced with BS |
30 Nov 2018 |
1.04 |
Day 04 - 4 Dec Raw Data |
DAY04 - LuTop, Mid and Bot postcal sysrsp - Lu rsp synced with BS |
5 Dec 2018 |
1.05 |
Day 05 - 5 Dec Raw Data |
DAY05 - Es wavecals |
5 Dec 2018 |
1.06 |
Day 06 - 6 Dec Raw Data |
DAY06 - LuTop/Bot wavecals |
6 Dec 2018 |
1.07 |
Day 07 - 7 Dec Raw Data |
DAY07 - LuTop int time cals |
10 Dec 2018 |
Daily data sets | |||
2.01 | check track defs | Check the track definitions for the Es and Lu tracks (looks good) | 4 Dec 2018 |
Day 2 data | |||
13.01 | Es system response data | Es system response data from Track 13 on day 3 using F601(c3) lamp file - need new wavecal to finalize | 4 Dec 2018 |
Day 4 data | |||
13.01 | Es/Lu system response | Es/Lu system response data from day 2 and 4 - need new wavecal to finalize | 5 Dec 2018 |
Day 5 data | |||
15.02 | Es wavlength cal | Es wavecals (HgA,Ne) | 5 Dec 2018 |
Day 6 data | |||
16.01 | LuTop wavlength cal | LuTop wavecals (all lamps) | 6 Dec 2018 |
16.02 | LuBot wavlength cal | LuBot wavecals (HgA,Ne) | 6 Dec 2018 |
Day 7 data - LuTop int time cal | |||
17.01 | int time cals | LuTop int time cals (Brite, Mid, dim and dimmer) | 10 Dec 2018 |
17.02 | int time cals 10 sec | LuTop int time cals (Brite, Mid, dim and dimmer) - normalized to 10 sec | 21 May 2019 |
17.03 | compare 1 and 10 sec cal | Comparing the 1 and 10 sec normalize int time cal | 22 May 2019 |
17.04 | SNR | SNR of the int time cal data | 29 May 2019 |
Compilation of day data | |||
20.01 | UV nets-> RS link | RS pow lix data are negative- We found some nevative values in the low pixel on H19-01 and are checking we see them here | 31 Jan 2019 |
DAY 1 EMAIL from Mike - LED data and Es rsps |
Hi Steph, |
DAY 4 EMAIL from Mike - Lu rsps |
Hi Steph, Last night was pos-MOBY265 cal = BS & RS Lu via OL425(L10). Zee data are at: /ftp1/Mike/L274/bs3cal /ftp1/Mike/L274/rs4cal It will be challenging to tune the responses of the BS & RS via iris'es and/or ND filters in the splitter, such that exposure times are compatible for Es & Ed & Lu over 2x spectrographs! For example, last night the exposure times (sec) : LuBot, BS=40, RS=3 LuMid, BS=43, RS=4.5 LuTop, BS=16, RS=16 So, last night I collected MANY more RS scan sets than BS scan sets for LuB & LuM, but the same number of scan sets for BS & RS via LuTop. I don't know that taking MANY more RS scan sets than BS scan sets was useful, however, I thought I'd do it this time to maybe look at the variability of the lamp via short-exp-time vs via long-exp-time - which might be more interesting in the water via sunlight vs tent + lamp! Q's: FileZilla continues to challenge me. It will not upload to the NAS for more than aboot 20 or 30 files before crashing. Then, sometimes, it will try to re-connect to the NAS and re-start uploads - which is never successful. So I have to Server/Disconnect the File/Exit and re-start FileZilla & figure out where to re-start uploads. This is the (repetative) error I get during file transfer failure: Response: 150 Opening ASCII mode data connection for d20181204079.fits Response: 425 Unable to build data connection: Operation not permitted Error: File transfer failed Q1: Have you ever tried / do you have trouble FTPing large numbers of files since you are now outside the MLML network/firewall kine? Q2: Do you know how to clear old "Queued files" from an old session? MF |
DAY 5 EMAIL from Mike - Es wavecals |
Hi Steph,
There are day5 data @ thine NAS for thee. Day5 twas just the HgA & Ne lamps via Es collector. I was not sure how many HgA lines would make it to the RS, so I added the Ne lamp just to be safe. The 1st Ne scan set was saturated @ the RS04, so I moved the lamp back from the Es diffuser & repeated both BS & RS scan sets via Ne. Note: BS03's Es is @ trk#13, and RS04's Es is @ trk#2 Tonight I plan to get all 5x wave lamps via LuTop = trk#6 on both BS & RS. Then tonight/tomorrow the Hga & Ne via LuBot = trk#2 & 13. Cheers, MF |
DAY 6 EMAIL from Mike - LuTop, Bot wavecals |
Hi Steph,
Day6 data are heading toward the NAS in starts & stops. They were via 5x wave lamps @ LuTop = trk#6 for both BS3 & RS4, plus HgA & Ne lamps @ LuBot = trk#2 for BS3 & trk#13 for RS4. Warning: all the Ne scans sets had saturated data... Luck, MF |
DAY 7 EMAIL from Mike - LuTop inttime cal |
Hi Steph,
Well, the VPN connection is SLOW, but it was steady = no errors! Thank you for your help setting this up. Day07 was exposure time cals for the BS & RS. The 2x filled-out log sheets for the BS apply to the RS (not filled out completely). The OL455 photodiode Amp output was logged to file = 2018120701_OL455log.txt, which is in the bs3cal dir, with the BS putty logs. Let me know if this looks OK! MF |